799,761,685,?,419,229
Answer by solver91311(24713)



799 - 761 = 38
761 - 685 = 76 = 2 X 38
685 - 419 = 266 = 7 X 38

So 419 + (4 X 38) = 571 and 571 + (3 X 38) = 685.

Your missing number is 571

John

My calculator said it, I believe it, that settles it